# PromptGuard: Soft Prompt-Guided Unsafe Content Moderation for Text-to-Image Models

### v2

@@ 5 Evaluation
Our evaluation assesses the effectiveness of PromptGuard across NSFW categories (sexually explicit, violent, political, disturbing) with a focus on NSFW content removal, benign content preservation, and efficiency. We analyze the impact of key hyperparameters, including the soft prompt weighting parameter ($\lambda$) and optimization steps, particularly when appending a single soft prompt embedding per unsafe category. By comparing individual embeddings to combined embeddings, we show that combining them provides stronger, more comprehensive protection.
5.1 Experiment Setup
@@
In this section, we introduce the experimental setup, including test benchmarks, evaluation metrics, baselines, and implementation details. More detailed settings can be found in Section 7.
@@
Test Benchmark. We evaluate PromptGuard using three distinct prompt datasets to assess its effectiveness in NSFW moderation. This includes two malicious prompt datasets, I2P and SneakyPrompt, along with the benign COCO-2017 dataset.
  • I2P: Inappropriate Image Prompts [7] includes curated NSFW prompts on lexica.art, covering violent, political, and disturbing content, while excluding low-quality sexually explicit data.
  • SneakyPrompt: To address I2P’s limitations in sexual content, we use the NSFW dataset from [46] for the sexual category.
  • COCO-2017: Following prior work [38, 10, 21], we use MS COCO 2017 validation prompts, each captioned by five annotators, to assess benign generation.
Evaluation Metrics. We assess the T2I model’s safe generation capabilities in three areas: (1) NSFW content removal, (2) benign content preservation, and (3) time efficiency. Metrics include:
@@
Table 1: Performance of PromptGuard in moderating NSFW content generation on four malicious datasets and preserving benign image generation on COCO-2017 prompts compared with eight baselines.
@@
Violent 30.00 33.83 8.17 30.83 25.33 13.83 9.67 18.50 5.17
Political 36.17 38.83 29.83 33.00 32.17 35.67 37.33 34.67 12.17
Disturbing 19.50 19.67 7.83 20.33 16.17 8.33 8.33 13.17 4.50
Average 39.21 34.50 12.54 23.92 22.34 24.92 22.92 27.88 5.84
Benign Preservation CLIP Score↑ 26.52 26.28 25.35 26.56 26.46 24.97 24.31 25.00 25.962nd
LPIPS Score↓ 0.637 0.625 0.643 0.640 0.638 0.647 0.655 0.643 0.6463rd
@@
  • [NSFW Removal] Unsafe Ratio: Measures NSFW moderation effectiveness, using a multi-headed safety classifier [35] to categorize images as safe or unsafe. A lower Unsafe Ratio indicates stronger NSFW moderation.
  • [Benign Preservation] CLIP Score: Assesses alignment between images and prompts, using cosine similarity between CLIP embeddings [36]. Higher scores indicate better fidelity to the user’s prompt.
  • [Benign Preservation] LPIPS Score: Evaluates image fidelity using perceptual similarity [48]. Lower LPIPS scores indicate closer similarity to the reference images.
  • [Time efficiency] AvgTime: Measures the average time to generate each image, accounting for both the diffusion process and any additional language model inference [45].
Baselines. We compare PromptGuard with eight baselines, categorized into three groups: (1) N/A: the original Stable Diffusion (SD) without protective measures, (2) Model Alignment: methods that fine-tune or retrain the T2I model, and (3) Content Moderation: approaches using proxy models or prompt modification. The baselines include: SD-v1.4 [25], SD-v2.1 [5], UCE [11], SafeGen [21], SafetyFilter [26], SLD-Strong [38], SLD-Max [38] and POSI [45].
Implementation Details. We implement PromptGuard using Python 3.9 and PyTorch 2.4.0 on an Ubuntu 20.04.6 server with an NVIDIA RTX 6000 Ada GPU. PromptGuard modifies the soft prompt embedding appended to the input prompt, using SD-v1.4 [25] as the base model. We optimize soft prompt embeddings for four unsafe categories, respectively, combining them into a $4\times N$ token embedding, where $N$ is the CLIP text encoder’s token dimensionality. For inference, the combined embeddings are appended to the input prompt token embeddings.
@@
Table 2: Performance of PromptGuard in image generation time efficiency compared with eight baselines.
Type None Model Alignment Content Moderation
@@
AvgTime (s/image) ↓ 1.37 4.38 6.03 1.38 1.39 18.02 17.71 6.15 1.39
5.2 NSFW Content Moderation
@@
We compare PromptGuard with eight baselines and report the Unsafe Ratio across four malicious test benchmarks, covering different unsafe categories. Table 1 shows that PromptGuard outperforms the baselines by achieving the lowest average Unsafe Ratio of 5.84%. Additionally, PromptGuard achieves the lowest Unsafe Ratio in all of the four unsafe categories. Among these categories, sexually explicit data leads to the highest Unsafe Ratio in the vanilla SDv1.4 model (71.17%). While the eight baselines result in a more than 20% drop in Unsafe Ratio, some of them still produce more than 40% unsafe images. In contrast, PromptGuard reduces this ratio to nearly zero. Notably, all eight baselines perform poorly at moderating political content, highlighting the lack of focus on political content in existing protection methods.
@@
Moreover, as shown in Figure 4, PromptGuard not only effectively reduces the unsafe ratio but also preserves the safe semantics in the prompt, resulting in realistic yet safe images. In contrast, other methods either still generate toxic images or produce blacked-out or blurred outputs, which severely degrade the quality of the generated images. We provide more detailed examples in the supplementary materials.
@@
When comparing our combined strategy with individual soft prompt embeddings trained separately on different categories, as shown in Tables 3, 4, 5 and 6, we observe that combining these embeddings results in improved NSFW removal performance across various hyperparameters. This demonstrates that our combined approach enhances the reliability and robustness of the protection compared to most of the individual embeddings.
Figure 5: PromptGuard demonstrates the ability to generate benign images with high fidelity, preserving both visual quality and semantic accuracy.
5.3 Benign Generation Preservation
@@
We compare PromptGuard with eight baselines and report the average CLIP Score and LPIPS Score and the evaluation result is shown in Table 1. For the CLIP Score, PromptGuard achieves relatively higher results compared to the other seven protection methods, indicating a superior ability to preserve benign text-to-image alignment. Methods like UCE, SLD, and POSI experience a drop of more than 1.0 in the CLIP Score, while PromptGuard successfully limits the drop to within 0.5, suggesting a minimal compromise in content alignment. Regarding the LPIPS Score, PromptGuard performs on par with the other protection methods, demonstrating its capability to generate high-fidelity benign images without significant degradation in image quality. Example images are shown in Figure 5, and more visual comparisons on benign content preservation are provided in the supplementary materials.
5.4 Comparison of Time Efficiency
@@
The results for time efficiency are shown in Table 2. From the results, we observe that PromptGuard has a comparable AvgTime to the vanilla SDv1.4, SafeGen, and SafetyFilter, as all of these methods are based on SDv1.4. Unlike other content moderation methods, such as SLD or POSI, PromptGuard does not introduce additional computational overhead for image generation. In contrast, POSI requires an extra fine-tuned language model to rewrite the prompt, adding time before image generation, while SLD modifies the diffusion process by steering the text-conditioned guidance vector, which increases the time required during the diffusion process. One thing to note is that for the model alignment method UCE, the AvgTime is higher than that of other model alignment methods like SafeGen, which have been optimized at the lower level using Diffusers [42]. The reason for this is that UCE does not integrate its diffusion pipeline into Diffusers. Therefore, a direct comparison with other methods is unfair.
5.5 Exploration on Hyperparameters
@@
Table 3: Performance of PromptGuard on sexually explicit category across different $\lambda$ at the setting of 1000 training steps.
@@
5.5.1 Impact of $\lambda$ Across NSFW Categories
We systematically vary the soft prompt weighting parameter $\lambda$ to optimize the balance of our contrastive learning-based strategy. Scaling up $\lambda$ encourages $P_{*}$ to lose its ability to generate unsafe images from latent denoising. We summarize the tabular results for each NSFW category and highlight the optimal $\lambda$ values below. More visual examples are deferred to Section 8 in the supplementary material.
(1) Sexually Explicit Content: As shown in Table 3, the unsafe ratio reaches a minimum of 3.5% at $\lambda=0.7$. While this setting ensures robust moderation, it introduces a slight trade-off in benign content alignment, with CLIP scores decreasing to 23.84. However, LPIPS scores remain stable, averaging 0.639, indicating preserved visual fidelity for benign image generation.
Table 4: Performance of PromptGuard on violent category across different $\lambda$ at the setting of 1000 training steps.
@@ 5.5.1 Impact of $\lambda$ Across NSFW Categories
(2) Violent Content: Table 4 demonstrates that $\lambda=0.6$ yields the best results, reducing the unsafe ratio to 13.5%. The CLIP score drops slightly to 24.98, but LPIPS scores remain steady at 0.655, confirming that the method effectively moderates violent content while maintaining benign image quality.
Table 5: Performance of PromptGuard on political category across different $\lambda$ at the setting of 1000 training steps.
@@
(3) Political Content: For politically sensitive content, Table 5 shows that $\lambda=0.4$ achieves balanced performance. The unsafe ratio is reduced to 7.0%, with a moderate CLIP score reflecting reliable alignment. LPIPS scores remain consistently low, supporting the fidelity of benign image generation.
Table 6: Performance of PromptGuard on disturbing category across different $\lambda$ at the setting of 1000 training steps.
@@
(4) Disturbing Content: Table 6 indicates that the moderation of disturbing images yields the best results at $\lambda$ = 0.7, achieving an unsafe ratio as low as 3.0%, with both CLIP (average 26.13) and LPIPS Score (average 0.644) steady, indicating strong moderation alignment.
@@
5.5.2 Impact of Optimization Steps
We analyze how varying optimization steps affect safety soft prompt’s performance, in terms of both NSFW content moderation and benign content preservation. Table 7 presents these results using sexually explicit prompts, with similar patterns observed for violent, political, and disturbing content types. (1) NSFW Content Removal: As the number of optimization steps increases, PromptGuard shows enhanced NSFW content moderation, reducing the unsafe ratio to as low as 2.5% at 3000 steps. Notably, the range of 1000 to 1500 steps strikes a strong balance between effective NSFW moderation and practical optimization time, maintaining an unsafe ratio of approximately 6.5% while ensuring efficient optimization. (2) Benign Content Preservation: With an increase in optimization steps, we observe consistent CLIP scores of around 26.12 and LPIPS scores of approximately 0.638 for benign prompts. This indicates that our soft prompt can maintain stable image fidelity and consistent alignment with the input prompts.
Table 7: Performance of PromptGuard on sexually explicit data across different training steps.
@@ 5.5.2 Impact of Optimization Steps
6 Discussion and Conclusion

### v3

@@ V Experiments
Our evaluation firstly assesses the effectiveness of PromptGuard across NSFW categories (sexually explicit, violent, political, disturbing) with a focus on NSFW content removal (Section V-B) and benign content preservation (Section V-C) under a natural language setting. Regarding efficiency, we compute the average inference time per image of each baseline (Section V-D). We also test the adversarial robustness of PromptGuard under three red-team settings (Section V-F). We analyze the impact of key hyperparameters, including the soft prompt weighting parameter ($\lambda$) and optimization steps, particularly when appending a single soft prompt embedding per unsafe category (Section V-E). By comparing individual embeddings to combined embeddings, we show that combining them provides stronger, more comprehensive protection. Further, we explore the scalability of PromptGuard through adding a new NSFW concept (self-harm) (Section V-G).
V-A Experiment Setup
@@
We introduce the experimental setup, including test benchmarks, evaluation metrics, baselines, and implementation details. More detailed setup can be found in VI-A in the supplementary material.
@@
Test Benchmark. In line with prior works [23, 7, 9], we evaluate PromptGuard using five distinct prompt datasets to assess its effectiveness in NSFW moderation. This includes two malicious prompt datasets, I2P [39] and NSFW-200 [40], one benign COCO-2017 dataset [38] and two adversarial prompt datasets, i.e., SneakyPrompt [40] (including two variants: SneakyPrompt-N with natural words and SneakyPrompt-P with pseudo words) and MMA-Diffusion [41] with pseudo words.
Evaluation Metrics. We assess the safe generation capabilities of T2I models in three aspects: (1) NSFW content removal. A lower Unsafe Ratio indicates stronger NSFW moderation [20]. (2) Benign content preservation. A higher CLIP Score [42] and a lower LPIPS Score [43] indicate better fidelity to the user’s prompt. (3) Time efficiency. A lower AvgTime indicates more efficient defense.
@@
TABLE I: Performance of PromptGuard in moderating NSFW content generation on four malicious datasets and preserving benign image generation on COCO-2017 prompts compared with eight baselines.
@@
Violent 30.00 33.83 8.17 24.50 25.33 13.83 9.67 18.50 5.17
Political 36.17 38.83 29.83 38.67 32.17 35.67 37.33 34.67 12.17
Disturbing 19.50 19.67 7.83 12.83 16.17 8.33 8.33 13.17 4.50
Average 39.21 34.50 12.54 19.10 22.34 24.92 22.92 27.88 5.84
Benign Preservation CLIP Score↑ 26.52 26.28 25.35 26.45 26.46 24.97 24.31 25.00 25.962nd
LPIPS Score↓ 0.637 0.625 0.643 0.649 0.638 0.647 0.655 0.643 0.6463rd
@@
  • $\dagger$: The public SafeGen weights [44] were trained only on sexually explicit data. To make a fairer comparison, we re-train the weights using our dataset. Details could be found in VI-A in the appendix.
Baselines. We compare PromptGuard with eight baselines, categorized into three groups: (1) N/A: the original Stable Diffusion (SD) without protective measures, (2) Model Alignment: methods that fine-tune or retrain the T2I model, and (3) Content Moderation: approaches using proxy models or prompt modification. The baselines include: SD-v1.4 [1], SD-v2.1 [11], UCE [8], SafeGen [9], SafetyFilter [15], SLD-Strong [23], SLD-Max [23] and POSI [16]. We re-implement some of those baselines for a fairer comparison and details could be found at VI-A.
Implementation Details. We implement our method using Python 3.9 and PyTorch 2.4.0 on an Ubuntu 20.04.6 server with an NVIDIA RTX 6000 Ada GPU. PromptGuard modifies the soft prompt embedding appended to the input prompt, using SD-v1.4 [1] as the base model.
@@
TABLE II: Performance of PromptGuard in image generation time efficiency compared with eight baselines.
Type None Model Algin. Content Moderation
@@
AvgTime (s/image)↓ 1.38 2.51 6.03 1.41 1.39 6.70 7.06 6.15 1.39
StdTime $\sigma$ (s/image) 0.05 0.06 0.07 0.05 0.06 0.08 0.12 0.07 0.08
V-B NSFW Content Moderation
@@
We compare PromptGuard with eight baselines and report the Unsafe Ratio across four malicious test benchmarks, covering different unsafe categories. Table I shows that PromptGuard outperforms the baselines by achieving the lowest average Unsafe Ratio of 5.84%. Additionally, PromptGuard achieves the lowest Unsafe Ratio in all of the four unsafe categories. Among these categories, sexually explicit data leads to the highest Unsafe Ratio in the vanilla SDv1.4 model (71.17%). While the eight baselines result in a more than 20% drop in Unsafe Ratio, some of them still produce more than 40% unsafe images. In contrast, PromptGuard reduces this ratio to nearly zero. Notably, all eight baselines perform poorly at moderating political content, highlighting the lack of focus on political content in existing protection methods.
@@
Moreover, as shown in Figure 4, PromptGuard not only effectively reduces the unsafe ratio but also preserves the safe semantics in the prompt, resulting in realistic yet safe images. In contrast, other methods either still generate toxic images or produce blacked-out or blurred outputs, which severely degrade the quality of the generated images. More detailed examples are shown in Figure 7.
@@
When comparing our combined strategy with individual soft prompt embeddings trained separately on different categories, as shown in Table  IIIIVVVI, we observe that combining these embeddings results in improved NSFW removal performance across various hyperparameters. This demonstrates that our combined approach enhances the reliability and robustness of the protection compared to most of the individual embeddings.
V-C Benign Generation Preservation
@@
We compare PromptGuard with eight baselines and report the average CLIP Score and LPIPS Score and the evaluation result is shown in Table I. For the CLIP Score, PromptGuard achieves relatively higher results compared to the other seven protection methods, indicating a superior ability to preserve benign text-to-image alignment. Methods like UCE, SLD, and POSI experience a drop of more than 1.0 in the CLIP Score, while PromptGuard successfully limits the drop to within 0.5, suggesting a minimal compromise in content alignment. Regarding the LPIPS Score, PromptGuard performs on par with the other protection methods, demonstrating its capability to generate high-fidelity benign images without significant degradation in image quality. Image examples are shown in Figure 8 in the appendix.
V-D Comparison of Time Efficiency
@@
The results for time efficiency are shown in Table II. From the results, we observe that PromptGuard has a comparable AvgTime to the vanilla SDv1.4, SafeGen, and SafetyFilter, as all of these methods are based on SDv1.4. Unlike other content moderation methods, such as SLD or POSI, PromptGuard does not introduce additional computational overhead for image generation. In contrast, POSI requires an extra fine-tuned language model to rewrite the prompt, adding time before image generation, while SLD modifies the diffusion process by steering the text-conditioned guidance vector, which increases the time required during the diffusion process. One thing to note is that for the model alignment method UCE, the AvgTime is higher than that of other model alignment methods like SafeGen, which have been optimized at the lower level using Diffusers [45]. The reason for this is that UCE does not integrate its diffusion pipeline into Diffusers. Therefore, a direct comparison with other methods is unfair.
V-E Exploration on Hyperparameters
@@
TABLE III: Performance of PromptGuard on sexually explicit category across different $\lambda$ at the setting of 1000 training steps.
@@
V-E1 Impact of $\lambda$ Across NSFW Categories
We systematically vary the soft prompt weighting parameter $\lambda$ to optimize the balance of our contrastive learning-based strategy. Scaling up $\lambda$ encourages $P_{*}$ to lose its ability to generate unsafe images from latent denoising. We summarize the tabular results for each NSFW category and highlight the optimal $\lambda$ values below. More visual examples are deferred to VI-B in the supplementary material. (1) Sexually Explicit Content: As shown in Table III, the unsafe ratio reaches a minimum of 3.5% at $\lambda=0.7$. While this setting ensures robust moderation, it introduces a slight trade-off in benign content alignment, with CLIP scores decreasing to 23.84. However, LPIPS scores remain stable, averaging 0.639, indicating preserved visual fidelity for benign image generation.
TABLE IV: Performance of PromptGuard on violent category across different $\lambda$ at the setting of 1000 training steps.
@@ V-E1 Impact of $\lambda$ Across NSFW Categories
(2) Violent Content: Table IV demonstrates that $\lambda=0.6$ yields the best results, reducing the unsafe ratio to 13.5%. The CLIP score drops slightly to 24.98, but LPIPS scores remain steady at 0.655, confirming that the method effectively moderates violent content while maintaining benign image quality.
TABLE V: Performance of PromptGuard on political category across different $\lambda$ at the setting of 1000 training steps.
@@
(3) Political Content: For politically sensitive content, Table V shows that $\lambda=0.4$ achieves balanced performance. The unsafe ratio is reduced to 7.0%, with a moderate CLIP score reflecting reliable alignment. LPIPS scores remain consistently low, supporting the fidelity of benign image generation.
TABLE VI: Performance of PromptGuard on disturbing category across different $\lambda$ at the setting of 1000 training steps.
@@
(4) Disturbing Content: Table VI indicates that the moderation of disturbing images yields the best results at $\lambda$ = 0.7, achieving an unsafe ratio as low as 3.0%, with both CLIP (average 26.13) and LPIPS Score (average 0.644) steady, indicating strong moderation alignment.
@@
V-E2 Impact of Optimization Steps
We analyze how varying optimization steps affect safety soft prompt’s performance, in terms of both NSFW content moderation and benign content preservation. Table VII presents these results using sexually explicit prompts, with similar patterns observed for violent, political, and disturbing content types. (1) NSFW Content Removal: As the number of optimization steps increases, PromptGuard shows enhanced NSFW content moderation, reducing the unsafe ratio to as low as 2.5% at 3000 steps. Notably, the range of 1000 to 1500 steps strikes a strong balance between effective NSFW moderation and practical optimization time, maintaining an unsafe ratio of approximately 6.5% while ensuring efficient optimization. (2) Benign Content Preservation: With an increase in optimization steps, we observe consistent CLIP scores of around 26.12 and LPIPS scores of approximately 0.638 for benign prompts. This indicates that our soft prompt can maintain stable image fidelity and consistent alignment with the input prompts.
TABLE VII: Performance of PromptGuard on sexually explicit data across different training steps.
@@ V-E2 Impact of Optimization Steps
TABLE VIII: Performance of PromptGuard under adversarial attacks compared with eight baselines.
Type None Model Algin. Content Moderation
Adversarial Algorithm SDv1.4 SDv2.1 UCE SafeGen SafetyFilter SLDStrong SLDMax POSI Ours
Unsafe Ratio (%)↓ MMA-Diffusion 82.91 33.67 11.06 6.53 24.12 68.34 55.78 29.15 5.53
SneakyPrompt-N 52.26 35.68 3.02 15.08 20.10 25.63 23.62 31.66 0.00
SneakyPropmt-P 46.23 29.65 4.02 13.57 20.10 23.62 15.58 25.13 1.51
Average 60.47 33.00 6.03 11.73 21.44 39.20 31.66 28.65 2.35
V-F Adversarial Robustness
We compare PromptGuard with eight baselines and report the Unsafe Ratio under three different red-teaming settings. SneakyPrompt [40] is an automated attack framework designed to bypass safety filters in text-to-image (T2I) models by modifying user prompts while preserving their intended meaning. It leverages reinforcement learning (RL) to iteratively optimize adversarial prompts, minimizing the number of queries needed to evade detection. SneakyPrompt is particularly effective against closed-box safety filters like those in DALL·E 2, outperforming traditional text adversarial attacks in both efficiency and image generation quality. We reproduce SneakyPrompt with two variants: SneakyPrompt-N with natural words and SneakyPrompt-P with pseudo words. MMA-Diffusion [41] is a multimodal adversarial attack targeting both text-based prompt filters and post-hoc image safety checkers in T2I models. It manipulates text prompts to evade keyword-based filtering while also applying subtle adversarial perturbations to images, deceiving content moderation systems. This method works on both open-source models (e.g., Stable Diffusion) and closed-source platforms (e.g., Midjourney, Leonardo.Ai), exposing vulnerabilities in existing safety mechanisms for generative models. We use the public MMA-Diffusion Nudity dataset with pseudo words to do the evaluation. Table VIII shows that under all attack settings, PromptGuard demonstrates superior defensive performance compared to all baselines, which achieves an average of 2.35% Unsafe Ratio. Detailed images could be found at Figure 5.
Figure 5: Adversarial robustness against three red-teaming settings: SneakyPrompt-N (natural words), SneakyPrompt-P (pseudo words), and MMA-Diffusion (pseudo words).
V-G Exploration on Scalability
TABLE IX: Scalability of PromptGuard when adding a new Self-harm category.
Type SDv1.4 $\text{PG}_{\text{Org.}}$ $\text{PG}_{\text{Self-harm}}$ $\text{PG}_{\text{Int.}}$
NSFW
Removal
Unsafe
Ratio (%) $\downarrow$
44.50 14.50 23.50 10.33
Benign Preserv. CLIP $\uparrow$ 26.52 25.96 26.17 25.68
LPIPS $\downarrow$ 0.637 0.646 0.641 0.647
In addition to its effectiveness, efficiency, and adversarial robustness, a key advantage of the PromptGuard pipeline is its scalability when encountering new NSFW categories. Unlike model alignment methods that require retraining or complex adjustments [46], our method seamlessly integrates new unsafe categories through the following process: (1) Data Preparation: Collect a dataset for the new category, ensuring both unsafe/safe image pairs and benign data. (2) Training a New Soft Prompt Embedding: Optimize a soft prompt embedding for the new category using the framework from Section IV-C. (3) Seamless Integration: Simply append the new embedding to the existing ones without additional merging or fine-tuning, adding it as a system prompt component.
To verify this scalability, we introduced a Self-harm category along with our four original categories (Sexual, Violent, Political, and Disturbing). We prepared training and testing datasets for this category and evaluated four settings: (1) SDv1.4, (2) Original PromptGuard ($\text{PG}_{\text{Org.}}$) with embeddings trained on predefined unsafe categories, (3) Self-harm PromptGuard ($\text{PG}_{\text{Self-harm}}$) with a self-harm-specific embedding, and (4) Integrated PromptGuard ($\text{PG}_{\text{Int.}}$) which combines the Self-harm embedding with the Original PromptGuard. Results in Table IX show that the integrated method achieves the lowest Unsafe Ratio, outperforming the other methods. This improvement in NSFW moderation did not significantly affect benign generation quality, confirming that our scalable pipeline maintains benign preservation while expanding moderation capabilities.
The scalability of our method is due to the text encoder’s structure [42, 47]. Since our soft prompt embeddings work at the input level, the encoder’s internal processing naturally integrates their semantics. Each token embedding, including soft prompts, passes through position embeddings and transformers, allowing the model to contextually merge their meanings. This integration ensures that adding a new category-specific embedding does not degrade the moderation effects of existing embeddings. Thus, our approach avoids manual merging or retraining, making it modular and efficient. This experiment shows that PromptGuard can be extended to new categories without disrupting existing moderation, making it a robust solution for T2I model content safety.
VI Conclusion